# Theory-agnostic searches for non-gravitational modes in black hole ringdown

### v1

@@ Abstract
Introduction. The black hole (BH) spectroscopy program [1, 2, 3, 4] plays a prominent role in the landscape of strong-field tests of General Relativity (GR) [5, 6, 7] and provides a unique method for examining the nature of compact remnants formed post-coalescence [8]. This program focuses on extracting the remnant quasinormal modes (QNMs) [9, 10, 11, 12] during the ringdown phase of a binary merger. In the context of linear perturbation theory, the signal $h(t)$ at intermediate times after the merger is represented by a superposition of the QNMs of the remnant [13], as it transitions towards a stationary configuration. Schematically,
@@
Standard tests of gravity based on the ringdown are rooted in the first line of Eq. (Theory-agnostic searches for non-gravitational modes in black hole ringdown). Namely, they are aimed at measuring $\delta f_{i}$ and $\delta\tau_{i}$ and check whether they are compatible with the null hypothesis. This can be accomplished in two complementary ways, either through a theory-agnostic or theory-dependent method. However, both approaches have their own limitations. In the theory-agnostic approach one uses the parametrization (2), where $f_{i}^{\rm Kerr}$ and $\tau_{i}^{\rm Kerr}$ are known functions of the BH mass and spin, while $\delta f_{i}$ and $\delta\tau_{i}$ depend on the mass, spin, and all extra fundamental coupling constants of the theory. It is convenient (and natural, from an effective-field-theory perspective) to restrict to small GR deviations that continuously deform the Kerr result, in which case $\delta f_{i},\delta\tau_{i}\ll 1$ and are proportional to some combination of the mass and coupling constants. However, even in this case they are still generic functions of the spin $\chi_{f}$. Current parametrizations either neglect such spin dependence [5, 34, 35] or consider a small spin expansion of each deviation [36, 37, 38], which inevitably inflates the number of free parameters in the model. In the theory-dependent approach, the QNMs are computed in a given theory of gravity. For most theories this can be done again only in the small-coupling limit and very often perturbatively in the spin [39, 40, 31, 41, 42, 43, 44, 45, 46, 47, 48]. To reach good convergence, one needs to push the spin expansion to very high order [47], which is very challenging from the technical point of view. Alternatively, one needs to solve intricate systems of coupled partial differential equations [49, 50, 51, 52, 53, 54]. This approach has the benefit of limiting the number of free parameters to the sole coupling constants and BH spin, but must be performed on a case-by-case basis for every given theory.
Given the above limitations, it would be highly desirable to develop complementary ringdown tests which are both theory-agnostic and accurate. In this work we explore a currently unbeaten path, related to the second line of Eq. (Theory-agnostic searches for non-gravitational modes in black hole ringdown). Namely, we propose to look for extra modes in the ringdown signal, which are not related to deformations of the Kerr ones. For concreteness, let us consider the case of an extra scalar degree of freedom nonminimally coupled to gravity (the same argument applies to other types of fields). Due to the coupling, the gravitational perturbations will contain also scalar modes (e.g., [40, 31, 55] for two concrete examples in theories with quadratic curvature terms),
@@
where $f_{i}^{{\rm Kerr},\,s=0}$ and $\tau_{i}^{{\rm Kerr},\,s=0}$ are the QNMs of a test scalar field in the Kerr metric, and also in this case $\delta\hat{f}_{i}$ and $\delta\hat{\tau}_{i}$ are complicated, theory-dependent, functions of the mass, spin, and coupling constants. Crucially, in this case the amplitudes $\hat{A}_{i}$ of these modes are proportional to (powers of) the coupling constants [40, 31, 55], and they must vanish in the GR limit. Therefore, to leading order in the corrections, we can neglect $\delta\hat{f}_{i}$ and $\delta\hat{\tau}_{i}$, so that the GR deviations are generically parametrized only by the amplitude of the test-field modes. This is precisely what happens in so-called dynamical Chern-Simons gravity [25, 31], although it is a generic feature.
@@
where for simplicity we have neglected the GR deviations in the first line, since those are very well studied by standard ringdown tests (and subjected to the aforementioned limitations). In practice, here we will focus on a standard GR ringdown waveform (first line of Eq. (Theory-agnostic searches for non-gravitational modes in black hole ringdown)) augmented by new extra modes. Remarkably, to leading order these extra modes are known functions of the BH mass and spin, since they are those of a free test (scalar, vector, etc) field propagating on the Kerr metric (see, e.g., [56, 57] for tabulated values). This allows searching for extra modes in a theory-agnostic way, where the amplitudes and phases of the extra modes are the only beyond-GR parameters. In this sense, this test is reminiscent of searches for extra (scalar, vector) polarizations in GW signals in a theory-agnostic fashion [58, 59, 60, 5] and is complementary to ordinary ringdown tests (see, e.g., [61, 5, 34, 35, 62, 63]).
Searching for extra ringdown modes. The ringdown signal comprises of two polarizations and the modes are decomposed in a basis of spin-weighted spheroidal harmonics that depend on the remnant spin inclination angle [63]. Each mode is identified by three integers, $i\equiv(lmn)$, namely the multipolar, azimuthal, and overtone index, respectively, where $n=0$ corresponds to the fundamental tone and $(lmn)=(220)$ is the dominant mode in a non-precessing quasicircular coalescence. For concreteness, here we focus on the most interesting quadrupolar ($l=2$) case, but our test can be applied also to higher-order modes. In particular, our ringdown waveform model has the following parameters:
@@
where $j={0,\dots,N}$, with $N$ the total number of overtones, whereas $\hat{A}_{220}^{s=0,1}$ and $\phi_{220}^{s=0,1}$ are the amplitude and phase of the extra scalar ($s=0$) or vector ($s=1$) $(220)$ mode. We will dub a model with $N$ tones and no extra mode GRN, whereas we will denote as GRN+S (GRN+V) a model with $N$ gravitational tones and an extra scalar (vector) mode. In practice, we define the amplitude ratio $A_{R}$ of a given mode relative to the gravitational fundamental one. We fix the luminosity distance $d_{L}$, the sky location of the system, and the inclination angle. Note that, for small-redshift sources (including those detected so far and presumably the loudest events detected in the future) $d_{L}$ is degenerate with the overall ringdown amplitude and can therefore be neglected without loss of generality. However, the possibility of neglecting the sky location and inclination angle is a prerogative of our test, since it involves only $(l=m=2)$ modes, which have the same pattern functions and spheroidal-harmonic decomposition. Thus, these extrinsic parameters factor out in the waveform, which is an important simplification compared to ordinary tests based on subleading modes with different values of $(l,m)$.
@@
Bayesian analysis on real data. We exemplify our test on real events by performing a Bayesian parameter estimation using the PyCBC Inference code infrastructure [64]. The analysis aims to compute the posterior distribution of the parameters (6). We apply this test to three events: (i) GW150914 [65], the first GW event ever detected by LIGO (and still so far the one with the largest ringdown signal-to-noise ratio (SNR)), for which some debated evidence of overtones has been reported [14, 66, 67, 68, 69, 70, 71, 72, 73, 34, 63]; (ii) GW190521 [74], a peculiar event in the upper mass gap for which a tentative detection of the $(330)$ has been obtained [75, 76] and which is prone also to ringdown amplitude-phase consistency tests [62]; (iii) GW200129, a peculiar loud event showing some tension with GR in some inspiral-merger-ringdown tests [5], tentatively ascribed to mismodelling of precession [77, 78]. For the luminosity distance, inclination, and sky location we adopt the maximum likelihood values reported in [79].
@@
Figure 2: Left panel: Minimum SNR necessary for detecting a scalar mode at $1-\sigma$ confidence level, according to the detectability criterion outlined in [82] using a Fisher matrix approximation. We take as a reference an event with remnant mass $M_{f}=62M_{\odot}$ for a range of amplitude ratios $A_{R,220}\in[0.02,1]$. The blue and red dashed curves denote a different remnant spin $\chi_{f}$ with the phase difference between the GR and scalar modes fixed to $\delta\phi=0$. The dependence on $\delta\phi\in[0,2\pi]$ is bracketed by the corresponding colored bands. The horizontal shaded bands represent the typical ringdown SNR expected for a GW150914-like event observed by the LVK network at design sensitivity (light blue), A+ (light green), ground-based third generation observatories as ET/CE (yellow), and LISA (orange). Finally, the blue dots denote the results of a Bayesian analysis obtained from injecting the scalar ringdown into zero noise for $\chi_{f}=0$. Right: The same quantity shown in a contour plot on the $(A_{R,220},\delta\phi)$ plane for fixed $\chi_{f}=0.67$.
@@
Forecasts with future observations. In Fig. 2, we forecast the constraining power of this test by computing the minimum ringdown SNR, $\rho_{\rm det}$, for detectability of an extra scalar mode in the GR0+S model, for different amplitude ratios, remnant spins, and phase differences $\delta\phi=\phi_{220}-\hat{\phi}_{220}^{s=0}$. (Hereafter we focus on the scalar case, since the vector case gives qualitatively similar results, see Appendix.) $\rho_{\rm det}$ is defined following Refs. [82, 84, 85], namely as the SNR such that $\sigma_{A_{R}}=A_{R,220}$, where the statistical error $\sigma_{A_{R}}$ has been computed with a fully numerical Fisher information matrix [84, 85]. Note that, in contrast to overtones [68, 82], resolving the scalar mode is relatively easy since the frequency is significantly different from the gravitational one. For this reason, $\rho_{\rm det}$ is always larger than the SNR threshold required to resolve the extra scalar mode.
@@
While these results were obtained using a Fisher-matrix approximation to explore the entire parameter space, in the left panel of Fig. 2 we also compare them with results obtained from synthetic injections at zero noise using the same Bayesian analysis discussed above for real-data events. The results of the full parameter estimation are in good agreement with the Fisher approximation, as expected given the large SNR.
Discussion. Any theory predicting extra ringdown modes should presumably predict also deviations from the standard gravitational Kerr QNMs, in which case one could argue that GW detectors are more sensitive to phase differences (and hence to QNM shifts) rather than amplitude differences, so that our test could have less constraining power than ordinary ringdown tests. However, there are known examples of theories predicting zero or negligible QNM shifts but extra modes, in which case our test can be superior to ordinary BH spectroscopy. An example is dynamical Chern-Simons gravity, wherein for a Schwarzschild BH the GR QNMs are unchanged, but the ordinary ringdown contains also extra scalar modes [31]. In the Kerr case also the GR QNMs are modified, but the deviations are suppressed by powers of the BH spin [44], so the convenience of our method will likely depend on the remnant’s spin.
@@
Acknowledgements. We would like to thank Collin Capano and Alexander Nitz for clarifications on PyCBC inference, and Gabriele Franciolini for insightful comments. This work is partially supported by the MUR PRIN Grant 2020KR4KN2 “String Theory as a bridge between Gauge Theories and Quantum Gravity”, by the FARE programme (GW-NEXT, CUP: B84I20000100001), and by the INFN TEONGRAV initiative. Some numerical computations have been performed at the Vera cluster supported by the Italian Ministry of Research and by Sapienza University of Rome, and with the University of Birmingham’s BlueBEAR HPC service.
Appendix A Supplemental results

### v2

@@ Abstract
Introduction. The black hole (BH) spectroscopy program [1, 2, 3, 4] plays a prominent role in the landscape of strong-field tests of General Relativity (GR) [5, 6, 7, 8] and provides a unique method for examining the nature of compact remnants formed post-coalescence [9]. This program focuses on extracting the remnant quasinormal modes (QNMs) [10, 11, 12, 13] during the ringdown phase of a binary merger. In the context of linear perturbation theory, the signal $h(t)$ at intermediate times after the merger is represented by a superposition of the QNMs of the remnant [14], as it transitions towards a stationary configuration. Schematically,
@@
Standard tests of gravity based on the ringdown are rooted in the first line of Eq. (Theory-agnostic searches for non-gravitational modes in black hole ringdown). Namely, they are aimed at measuring $\delta f_{i}$ and $\delta\tau_{i}$ and check whether they are compatible with the null hypothesis [5]. This can be accomplished in two complementary ways, either through a theory-agnostic or theory-dependent method. However, both approaches have their own limitations. In the theory-agnostic approach one uses the parametrization (2), where $f_{i}^{\rm Kerr}$ and $\tau_{i}^{\rm Kerr}$ are known functions of the BH mass and spin, while $\delta f_{i}$ and $\delta\tau_{i}$ depend on the mass, spin, and all extra fundamental coupling constants of the theory. It is convenient (and natural, from an effective-field-theory perspective) to restrict to small GR deviations that continuously deform the Kerr result, in which case $\delta f_{i},\delta\tau_{i}\ll 1$ and are proportional to some combination of the mass and coupling constants. However, even in this case they are still generic functions of the spin $\chi_{f}$. Current parametrizations either neglect such spin dependence [5, 35, 36] or consider a small spin expansion of each deviation [37, 38, 39], which inevitably inflates the number of free parameters in the model. In the theory-dependent approach, the QNMs are computed in a given theory of gravity. For most theories this can be done again only in the small-coupling limit and very often perturbatively in the spin [40, 41, 32, 42, 43, 44, 45, 46, 47, 48, 49]. To reach good convergence, one needs to push the spin expansion to very high order [48], which is very challenging from the technical point of view. Alternatively, one needs to solve intricate systems of coupled partial differential equations [50, 51, 52, 53, 54, 55]. This approach has the benefit of limiting the number of free parameters to the sole coupling constants and BH spin, but must be performed on a case-by-case basis for every given theory.
Given the above limitations, it would be highly desirable to develop complementary ringdown tests which are both theory-agnostic and accurate. In this work we explore a currently unbeaten path, related to the second line of Eq. (Theory-agnostic searches for non-gravitational modes in black hole ringdown). Namely, we propose to look for extra modes in the ringdown signal, which are not related to deformations of the Kerr ones. For concreteness, let us consider the case of an extra scalar degree of freedom nonminimally coupled to gravity (the same argument applies to other types of fields). Due to the coupling, the gravitational perturbations will contain also scalar modes (e.g., [41, 32, 56, 34] for two concrete examples in theories with quadratic curvature terms),
@@
where $f_{i}^{{\rm Kerr},\,s=0}$ and $\tau_{i}^{{\rm Kerr},\,s=0}$ are the QNMs of a test scalar field in the Kerr metric, and also in this case $\delta\hat{f}_{i}$ and $\delta\hat{\tau}_{i}$ are complicated, theory-dependent, functions of the mass, spin, and coupling constants. Crucially, in this case the amplitudes $\hat{A}_{i}$ of these modes are proportional to (powers of) the coupling constants [41, 32, 56, 34], and they must vanish in the GR limit. Therefore, to leading order in the corrections, we can neglect $\delta\hat{f}_{i}$ and $\delta\hat{\tau}_{i}$, so that the GR deviations are generically parametrized only by the amplitude of the test-field modes. This is precisely what happens in so-called dynamical Chern-Simons gravity [26, 32, 34] (see also Appendix), although it is a generic feature.
@@
where for simplicity we have neglected the GR deviations in the first line, since those are very well studied by standard ringdown tests (and subjected to the aforementioned limitations). In practice, here we will focus on a standard GR ringdown waveform (first line of Eq. (Theory-agnostic searches for non-gravitational modes in black hole ringdown)) augmented by new extra modes. Remarkably, to leading order these extra modes are known functions of the BH mass and spin, since they are those of a free test (scalar, vector, etc) field propagating on the Kerr metric (see, e.g., [57, 58] for tabulated values). This allows searching for extra modes in a theory-agnostic way, where the amplitudes and phases of the extra modes are the only beyond-GR parameters. In this sense, this test is reminiscent of searches for extra (scalar, vector) polarizations in GW signals in a theory-agnostic fashion [59, 60, 61, 5] and is complementary to ordinary ringdown tests (see, e.g., [62, 63, 5, 35, 36, 64, 65]), or to test with multiple free modes [5].
Searching for extra ringdown modes. The ringdown signal comprises of two polarizations and the modes are decomposed in a basis of spin-weighted spheroidal harmonics that depend on the remnant spin inclination angle [65]. Each mode is identified by three integers, $i\equiv(lmn)$, namely the multipolar, azimuthal, and overtone index, respectively, where $n=0$ corresponds to the fundamental tone and $(lmn)=(220)$ is the dominant mode in a non-precessing quasicircular coalescence. For concreteness, here we focus on the most interesting quadrupolar ($l=2$) case and neglect spin precession of the progenitor binary, but our test can be applied also to higher-order modes and to the precessing case (see Appendix). In particular, our ringdown waveform model has the following parameters:
@@
where $j={0,\dots,N}$, with $N$ the total number of overtones, whereas $\hat{A}_{220}^{s=0,1}$ and $\phi_{220}^{s=0,1}$ are the amplitude and phase of the extra scalar ($s=0$) or vector ($s=1$) $(220)$ mode. We will dub a model with $N$ tones and no extra mode GRN, whereas we will denote as GRN+S (GRN+V) a model with $N$ gravitational tones and an extra scalar (vector) mode. In practice, we define the amplitude ratio $A_{R}$ of a given mode relative to the gravitational fundamental one. We fix the luminosity distance $d_{L}$, the sky location of the system, and the inclination angle. Note that, for small-redshift sources (including those detected so far and presumably the loudest events detected in the future) $d_{L}$ is degenerate with the overall ringdown amplitude and can therefore be neglected without loss of generality. However, the possibility of neglecting inclination angle is a prerogative of our test, since it involves only $(l=m=2)$ modes, which have the same pattern functions and spheroidal-harmonic decomposition. Compared to common tests based on subleading modes of different $(l,m)$, this allows for the practical simplification of sampling in only one of these three degenerate parameters.
@@
Bayesian analysis on real data. We exemplify our test on real events by performing a Bayesian parameter estimation using the PyCBC Inference code infrastructure [66]. The analysis aims to compute the posterior distribution of the parameters (6). We apply this test to three events: (i) GW150914 [67], the first GW event ever detected by LIGO (and still so far the one with the largest ringdown signal-to-noise ratio (SNR)), for which some debated evidence of overtones has been reported [15, 68, 69, 70, 71, 72, 73, 74, 75, 76, 35, 65]; (ii) GW190521 [77], a peculiar event in the upper mass gap for which a tentative detection of the $(330)$ and other modes (and possibly precession) has been obtained [78, 79] and which is prone also to ringdown amplitude-phase consistency tests [64]; (iii) GW200129, a peculiar loud event showing some tension with GR in some inspiral-merger-ringdown tests [5], tentatively ascribed to mismodelling of precession [80, 81]. As a proof of principle for the test, our parametrization assumes plane reflection symmetry, which is valid for spin-aligned progenitor binaries. In the Appendix we show that relaxing this assumption does not affect the results for GW190521 and GW200129. For the luminosity distance, inclination, and sky location we adopt the maximum likelihood values reported in [82].
@@
Figure 2: Left panel: Minimum SNR necessary for detecting a scalar mode at $1-\sigma$ confidence level, according to the detectability criterion outlined in [85] using a Fisher matrix approximation. We take as a reference an event with remnant mass $M_{f}=62M_{\odot}$ for a range of amplitude ratios $A_{R,220}\in[0.02,1]$. The blue and red dashed curves denote a different remnant spin $\chi_{f}$ with the phase difference between the GR and scalar modes fixed to $\delta\phi=0$. The dependence on $\delta\phi\in[0,2\pi]$ is bracketed by the corresponding colored bands. The horizontal shaded bands represent the typical ringdown SNR expected for a GW150914-like event observed by the LVK network at design sensitivity (light blue), A+ (light green), ground-based third-generation observatories as ET/CE (yellow), and for a mass-rescaled version of GW150914 with $10^{5}M_{\odot}$ for LISA (orange). Right: The same quantity shown in a contour plot on the $(A_{R,220},\delta\phi)$ plane for fixed $\chi_{f}=0.67$.
@@
Forecasts with future observations. In Fig. 2, we forecast the constraining power of this test by computing the minimum ringdown SNR, $\rho_{\rm det}$, for detectability of an extra scalar mode in the GR0+S model, for different amplitude ratios, remnant spins, and phase differences $\delta\phi=\phi_{220}-\hat{\phi}_{220}^{s=0}$. (Hereafter we focus on the scalar case, since the vector case gives qualitatively similar results, see Appendix.) $\rho_{\rm det}$ is defined following Refs. [85, 87, 88], namely as the SNR such that $\sigma_{A_{R}}=A_{R,220}$, where the statistical error $\sigma_{A_{R}}$ has been computed with a fully numerical Fisher information matrix [87, 88]. Note that, in contrast to overtones [70, 85], resolving the scalar mode is relatively easy since the frequency is significantly different from the gravitational one. For this reason, $\rho_{\rm det}$ is always larger than the SNR threshold required to resolve the extra scalar mode. We perform two variants of this analysis, either averaging over or fixing the sky-location of the signal, with results shown in the left and right panel of Fig. 2, respectively.
@@
While these results were obtained using a Fisher-matrix approximation to explore the entire parameter space, we have also compared individual points with synthetic injections at zero noise using the same Bayesian analysis discussed above for real-data events. At SNRs of $1.25\rho_{\text{det}}$ as given in Fig. 2, we find posteriors of $A_{R}$ peaking away from the lower bound and $\sigma_{A_{R}}\lesssim A_{R,220}$, already in agreement with the expectation from the Fisher analysis for the high-SNR limit.
Discussion. Any theory predicting extra ringdown modes should presumably also predict deviations from the standard gravitational Kerr QNMs, in which case one could argue that GW detectors are more sensitive to phase differences (and hence to QNM shifts) rather than amplitude differences, so that our test could have less constraining power than ordinary ringdown tests. However, there are known examples of theories predicting zero or negligible QNM shifts but extra modes, in which case our test can be superior to ordinary BH spectroscopy. An example is dynamical Chern-Simons gravity, wherein for a Schwarzschild BH the GR QNMs are unchanged, but the ordinary ringdown contains also extra scalar modes [32]. In the Kerr case also the GR QNMs are modified, but the deviations are suppressed by powers of the BH spin [45], so the convenience of our method will likely depend on the remnant’s spin.
@@
Acknowledgements. We thank Collin Capano and Alexander Nitz for clarifications on PyCBC inference, and Gregorio Carullo, Cecilia Chirenti, Gabriele Franciolini, Nico Yunes, and the participants of the workshop Ringdown Inside and Out for insightful comments and discussions. This work is partially supported by the MUR PRIN Grant 2020KR4KN2 “String Theory as a bridge between Gauge Theories and Quantum Gravity”, by the FARE programme (GW-NEXT, CUP: B84I20000100001), and by the INFN TEONGRAV initiative. X. Jimenez is supported by the Spanish Ministerio de Ciencia, Innovación y Universidades (Beatriz Galindo, BG22-00034) and cofinanced by UIB; the Spanish Agencia Estatal de Investigación Grants No. PID2022-138626NB-I00, No. RED2022- 134204-E, and No. RED2022-134411-T, funded by MCIN/AEI/10.13039/501100011033/FEDER, UE; the MCIN with funding from the European Union NextGenerationEU/PRTR (No. PRTR-C17.I1); the Comunitat Autonòma de les Illes Balears through the Direcció General de Recerca, Innovació I Transformació Digital with funds from the Tourist Stay Tax Law (No. PDR2020/11 - ITS2017-006), and the Conselleria d’Economia, Hisenda i Innovació Grant No. SINCO2022/6719. Some numerical computations have been performed at the Vera cluster supported by the Italian Ministry of Research and by Sapienza University of Rome, and with the University of Birmingham’s BlueBEAR HPC service.
Appendix A Supplemental material
A.1 Explicit example: Dynamical Chern-Simons gravity
We provide here a specific example of a nonminimal coupling giving rise to extra scalar modes in the gravitational sector. We consider a theory with quadratic curvature corrections, dynamical Chern-Simons gravity, described by the action [26]
$S=\frac{1}{16\pi}\int d^{4}x\sqrt{-g}R-\frac{1}{2}\int d^{4}x\sqrt{-g}g^{ab}\nabla_{a}\vartheta\nabla_{b}\vartheta$
$+\frac{\alpha}{4}\int d^{4}x\sqrt{-g}\vartheta\,^{*}RR\,.$ (8)
where $\vartheta$ is the scalar field, ${}^{*}RR=\frac{1}{2}R_{abcd}\epsilon^{baef}R^{cd}_{~{}~{}ef}$ is an odd-parity quadratic-curvature invariant, and $\alpha$ is the coupling constant, with dimensions of a squared mass (henceforth we adopt $G=c=1$ units).
As in the GR case, the only stationary, spherically-symmetric solution is the Schwarzschild metric. For simplicity we consider perturbations of this solution, neglecting the spin of the background. Axial perturbations of the metric are coupled to those of the scalar field. Upon a spherical harmonic decomposition and in the frequency domain, they reduce to the following set of coupled ordinary differential equations [32]
$\frac{d^{2}}{dr_{\star}^{2}}\Psi+\left\{\omega^{2}-f\left[\frac{l(l+1)}{r^{2}}-\frac{6M}{r^{3}}\right]\right\}\Psi$ $=$ $\frac{96\pi\,Mf}{r^{5}}\alpha\Theta\,,$ (9)
$\frac{d^{2}}{dr_{\star}^{2}}\Theta+\left\{\omega^{2}-f\left[\frac{l(l+1)}{r^{2}}\left(1+\frac{576\pi M^{2}\alpha^{2}}{r^{6}}\right)+\frac{2M}{r^{3}}\right]\right\}\Theta$ $=$ $f\frac{(l+2)!}{(l-2)!}\frac{6M\alpha}{r^{5}}\Psi$ (10)
where $f(r)=1-2M/r$ and $r_{\star}\equiv r+2M\ln\left(r/2M-1\right)$ is the standard Schwarzschild tortoise coordinate. The variables $\Psi$ and $\Theta$ reduce to the standard metric and scalar master functions, respectively, in the decoupling limit, $\alpha\to 0$. Indeed, when $\alpha=0$ the two equations decouple and reduce to the standard Regge-Wheeler equation and scalar-perturbation equation of a Schwarzschild BH, respectively. However, when $\alpha\neq 0$ the two perturbations are coupled to each other and the scalar effective potential acquires some corrections.
The coupling $\alpha$ gives rise to two features [32, 34]:
  1. 1.
    The above system of equations contains both gravity-led and scalar-led modes, both displaying ${\cal O}(\alpha^{2})$ corrections with respect to GR:
    $\omega$ $=\omega^{\rm GR,~{}grav}\left(1+\gamma^{\rm grav}\frac{\alpha^{2}}{M^{4}}\right)\,,$ (11)
    $\hat{\omega}$ $=\omega^{\rm GR,~{}scal}\left(1+\gamma^{\rm scal}\frac{\alpha^{2}}{M^{4}}\right)\,,$ (12)
    where $\omega=2\pi f-i/\tau$ are the complex QNM frequencies, $\omega^{\rm GR,~{}grav}$ are the standard Schwarzschild QNMs in GR, $\omega^{\rm GR,~{}scal}$ are the test-scalar QNMs of Schwarzschild, whereas $\gamma^{\rm grav}$ and $\gamma^{\rm scal}$ are dimensionless order-unity constants, the value of which depends on the overtone number $n$.
  2. 2.
    The above system of equations is akin to a coupled harmonic oscillator, so a scalar perturbation would source scalar modes in the gravitational sector, and vice versa. In particular, in this theory the scalar field is at least linear in $\alpha$ [26], in which case the coupled perturbation equations imply that, whether or not scalar perturbations are present in the merger conditions (for example if the progenitors are endowed with a scalar field [106]), the amplitude of the scalar mode in the gravitational sector would be ${\cal O}(\alpha^{2})$. In the notation of the main text, this would imply that, at the leading order, $A_{R,220}\propto\alpha^{2}/M_{f}^{4}$, which is the same order of the usual leading-order corrections to the QNMs in this theory [32, 45, 49].
These arguments show that the gravitational ringdown in this theory can be schematically modelled as
$h(t)=$ $\sum_{j}A_{j}e^{i(\omega_{j}t+\phi_{j})}+\sum_{j}\hat{A}_{j}e^{i(\hat{\omega}_{j}t+\hat{\phi}_{j})}\,,$ (13)
where the sum runs over the overtones. The scalar modes $\hat{\omega}_{j}$ contains ${\cal O}(\alpha^{2})$ corrections, but $\hat{A}_{j}$ is at least ${\cal O}(\alpha)$ or higher. Therefore, to leading order in $\alpha$ we can approximate $\hat{\omega}_{j}\approx\omega_{j}^{\rm GR,~{}scal}$ in the second term of the above equation, which is the crucial simplification of our test. This is consistent with the finding of Ref. [32], where the gravitational ringdown in the $\alpha\ll M^{2}$ limit contains both the unperturbed gravitational and scalar modes.
While we explicitly showed this for a specific theory, it is in fact a very general properties of extended theories of gravity with nonminimal couplings (see, e.g., [56] for another example).
A.2 The effect of precession on GW190521 and GW200129
Figure 3: Posterior distributions for the event GW200129 with $t_{\text{offset}}=0\text{ms}$ for the models GR1 (left) and GR0+S (right). In both cases, the case without (resp., with) precession is indicated in yellow (resp., purple).
There are well-founded arguments suggesting that GW190521 and GW200129 could originate from precessing binaries [81, 79]. As opposed to aligned-spin systems, precession breaks the symmetry between $m\leftrightarrow-m$ angular modes, resulting in $h_{lm}\neq(-1)^{l}h^{*}_{l-m}$. In the main text, we searched for a scalar mode using a $(lm)=(2,\pm 2)$ spin-aligned (non-precessing) waveform. Here, we extend the parameter estimation on GW190521 and GW200129 including precession, and varying $t_{\text{offset}}=[-2,0,2]\text{ms}$. This is achieved by allowing the amplitude $A_{22n}\neq A_{2-2n}$ and the phase $\phi_{22n}\neq-\phi_{2-2n}$, for both the gravitational and scalar mode. In Fig. 3, we show the corner plots for GW200129 with $t_{\text{offset}}=0$, for both the GR01 and GR0+S cases. Notice that, since only $(lm)=(2,\pm 2)$ modes are used, accounting for precession in this case does not significantly affect the values of the mass and the spin of the remnant compared to the aligned-spin case. However, including precession affects the amplitude of the fundamental mode, by increasing its uncertainty. We find qualitatively similar results for the other offset times and for GW190521.
A.3 Supplemental results